# thanks: These authors contribute equally to this workthanks: These authors contribute equally to this workthanks: These authors contribute equally to this work

### v2

@@ Abstract
Acknowledgements This work is supported by Innovation Program for Quantum Science and Technology 2021ZD0302005 (H.Z.), 2024ZD0300101 (P.Z.), and 2021ZD0303205 (X.P.), National Natural Science Foundation of China No. 12488301 (H.Z.), No. U23A6004 (H.Z.), No. 12374477 (P.Z.), No. 12261160569 (X.P.), the Shanghai Rising-Star Program under grant number 24QA2700300 (P.Z.), and the XPLORER Prize (H.Z. and X.P.). We thank experimental assistance from the Instruments Center for Physical Science, University of Science and Technology of China.
Data and Code Availability The raw NMR experimental data and all codes are available at [50].

### v3

@@ Abstract
Acknowledgements—This work is supported by Quantum Science and Technology-National Science and Technology Major Project 2021ZD0302005 (H.Z.), 2024ZD0300101 (P.Z.), and 2021ZD0303205 (X.P.), National Natural Science Foundation of China No. 12488301 (H.Z.), No. U23A6004 (H.Z.), No. 12374477 (P.Z.), No. 12261160569 (X.P.), the Shanghai Rising-Star Program under grant number 24QA2700300 (P.Z.), and the XPLORER Prize (H.Z. and X.P.). We thank experimental assistance from the Instruments Center for Physical Science, University of Science and Technology of China.
Data Availability—The data that support the findings of this study are available at Zenodo (https://doi.org/10.5281/zenodo.16142455) [70]. Further data are available from the corresponding authors upon reasonable request.